Charlie owns a coffee shop. He wants to create special mixes using 2 coffees,one priced at $6.40 per pound and the other priced at $7.28 per pound. How many pounds of the $7.28 coffee should he mix with 9 pounds of the $6.40 coffee to sell the mixture for $6.95 per pound?
Answer by nerdybill(5412)

Charlie owns a coffee shop. He wants to create special mixes using 2 coffees,one priced at $6.40 per pound and the other priced at $7.28 per pound. How many pounds of the $7.28 coffee should he mix with 9 pounds of the $6.40 coffee to sell the mixture for $6.95 per pound?
.
Let x = pounds of $7.28 per pound coffee added
.
7.28x + 6.40(9) = 6.95(x+9)
7.28x + 57.6 = 6.95x + 62.55
0.33x + 57.6 = 62.55
0.33x = 4.95
x = 15 pounds (of $7.28 coffee)